Dog Up! Franks is looking at a new sausage system with an installed cost of $312,000. This cost will be depreciated straight-line to zero over the project's 4-year life, at the end of which the sausage system can be scrapped for $48,000. The sausage system will save the firm $96,000 per year in pretax operating costs, and the system requires an initial investment in net working capital of $22,400. If the tax rate is 33 percent and the discount rate is 16 percent, what is the NPV of this project?
Answer #1

rate positively..

Year 0 1 2 3 4
A initial investment -312000
B working capital -22400 22400
Operating cash flow
i Cost saving 96000 96000 96000 96000
ii Depreciation 78000 78000 78000 78000
iii Net saving 18000 18000 18000 18000
iv Tax @ 33% 5940 5940 5940 5940
v Profit before tax 12060 12060 12060 12060
C=v+ii Cash flow 90060 90060 90060 90060
Terminal value =
D 48000*(1-33%) 32160
E=A+B+C+D Net cash flow -334400 90060 90060 90060 144620
F PVIF @ 16% 1 0.862069 0.7431629 0.6406577 0.5522911
G=E*F present value (334,400.00) 77,637.93 66,929.25 57,697.63 79,872.34 (52,262.85)
NPV =     (52,262.85)